Find the monthly PITI payment for a mortgage of $196,000 at a rate of 6.5% for 25 years with annual property taxes and hazard insurance premium of $4,850 and $1084 respectively.
Answer by amarjeeth123(569)

The monthly PITI payment for the mortgage of 196,000atarateof6.5 4,850 and 1,084respectivelyis 1,341.60.
How to calculate the amount
We need to use the following formula:
PMT = (Pr(1+r)^n)/((1+r)^n-1) + T + I
Where:
PMT is the monthly payment
Using this formula, we can calculate the monthly PITI payment for the given mortgage as follows:
P = 196,000
r = 6.5% / 12 = 0.00541667
n = 25 * 12 = 300
T = 4,850 / 12 = 404.17
I = 1,084 / 12 = 90.33
PMT = (1960000.00541667(1+0.00541667)^300)/((1+0.00541667)^300-1) + 404.17 + 90.33
PMT = 1,341.6